# Reinforcement Learning for Sequence Design Leveraging Protein Language Models

### v1

@@ 3 Preliminaries
We frame the problem of protein sequence design in an RL loop, where at each timestep the agent receives a sequence and predicts the mutation to be performed. The sequences are represented through one hot encodings of each character in the sequence. That is, given a sequence of length $L_{s}$ and an alphabet of size $L_{a}$, the state vector is denoted by $s\in\mathbb{R}^{L_{s}\times L_{a}}$. The action space is modelled as a discrete choice over $L_{s}\times L_{a}$ options, indicating the position of the mutation and the mutated character. The mutated sequence is then returned as the next state. The reward is the pTM score of the mutated sequence, which is obtained from either the oracle model or the proxy model. Our environment thus is completely deterministic, the state transitions are simple sequence mutations and the rewards are deterministically obtained from the models. Finally, we design the environment to be an infinite horizon task. Since querying the reward model is expensive, we create multiple environments in parallel and query the model with a batch of states.
@@ 5 Experiments and Analysis
Table 1 and 4 (appendix) present the results for all methods evaluated in the infinite horizon setting for sequences of lengths $50$ and $100$, respectively. The reward used in these experiments is the pTM scores obtained from the oracle model (ESMFold). All results are averaged over three random seeds, and all experiments for this analysis were run on 4 A100 GPUs. The results in these tables highlight the effectiveness of RL algorithms and GFlowNets for sequence design.
@@
We find that all RL methods and GFlowNets are able to maintain high oracle pTM scores by optimizing on the proxy pTM (Fig  4 and Table 3), while cutting down the runtime by an order of magnitude (Table  5). All results are averaged over three random seeds, and all experiments for this analysis were run on a single RTX8000 GPU. We present the complete results in Table  3.
@@
Although MCMC is effective for direct optimization, it struggles to generate sequences effectively when used in conjunction with a continually finetuned proxy. This limitation becomes particularly significant when computational resources are constrained. In contrast, other methods show greater robustness when training with the proxy, indicating their potential for more reliable performance under such conditions.
@@ Acknowledgements
The authors would like to thank Compute Canada and Mila for the computing resources required for this work. JS would like to thank Kausthubh Ramachandran, Michael Jendrusch, Moksh Jain and Nanda Harishankar Krishna for fruitful discussions and for providing feedback on initial drafts of the paper.
@@ 75. Guidelines:
Appendix A Appendix / supplemental material
Figure 7: Pareto plots for optimization on ESMFold scores: Plotted across two structural diversity metrics and one sequence diversity metric (MP-HD). Only methods with atleast $0.5$ pTM are highlighted.
@@ Appendix A Appendix / supplemental material
Protein Confidence Diversity Score pTM ($\uparrow$) pLDDT ($\uparrow$) MP-TM ($\downarrow$) MP-RMSD ($\uparrow$) MP-HD ($\uparrow$) Random 0.40$\pm$0.00 46.9$\pm$0.48 0.35$\pm$0.01 3.99$\pm$0.02 0.89$\pm$0.02 MH-MCMC 0.93$\pm$0.00 92.2$\pm$1.09 1.00$\pm$0.00 0.04$\pm$0.03 0.01$\pm$0.01 DQN 0.37$\pm$0.03 42.2$\pm$2.74 0.62$\pm$0.14 2.76$\pm$0.75 0.15$\pm$0.12 SAC 0.71$\pm$0.02 78.5$\pm$2.32 0.71$\pm$0.12 1.87$\pm$0.43 0.66$\pm$0.06 PPO 0.86$\pm$0.02 88.9$\pm$1.90 1.00$\pm$0.00 0.03$\pm$0.03 0.01$\pm$0.01 PPO-RND 0.50$\pm$0.05 55.4$\pm$5.16 0.91$\pm$0.11 0.53$\pm$0.64 0.16$\pm$0.20 Rainbow 0.45$\pm$0.06 52.7$\pm$3.70 1.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 GFN $(\beta=6)$ 0.42$\pm$0.03 48.6$\pm$1.13 0.35$\pm$0.01 3.99$\pm$0.11 0.90$\pm$0.01 GFN $(\beta=32)$ 0.46$\pm$0.01 53.7$\pm$1.39 0.38$\pm$0.01 3.59$\pm$0.10 0.88$\pm$0.02 GFN $(\beta=64)$ 0.55$\pm$0.04 63.6$\pm$1.87 0.55$\pm$0.03 2.66$\pm$0.33 0.78$\pm$0.07

### v2

@@ 3 Preliminaries
We frame the problem of protein sequence design in an RL loop, where at each timestep the agent receives a sequence and predicts the mutation to be performed. The sequences are represented through one hot encodings of each character in the sequence. That is, given a sequence of length $L_{s}$ and an alphabet of size $L_{a}$, the state vector is denoted by $s\in\mathbb{R}^{L_{s}\times L_{a}}$. The action space is modelled as a discrete choice over $L_{s}\times L_{a}$ options, indicating the position of the mutation and the mutated character. The mutated sequence is then returned as the next state. The reward is the pTM score of the mutated sequence, which is obtained from either the oracle model or the proxy model. Our environment thus is completely deterministic, the state transitions are simple sequence mutations and the rewards are deterministically obtained from the models. Finally, we design the environment to be an infinite horizon task. RL training is simultaneously conducted on a batch of $100$ sequences. ESMFold is queried with the batch of all $100$ sequences and inference is run across 4 GPUs to obtain the rewards (i.e, pTM score).
@@ 5 Experiments and Analysis
Table 1 and 4 (appendix) present the results for all methods evaluated in the infinite horizon setting for sequences of lengths $50$ and $100$, respectively. The reward used in these experiments is the pTM scores obtained from the oracle model (ESMFold). All results are averaged over three random seeds, and all experiments for this analysis were run on 4 A100 GPUs. The results in these tables highlight the effectiveness of RL algorithms and GFlowNets for sequence design. Figure 3 illustrates the training curves for optimizing the oracle reward for sequences of length $50$.
@@
We find that all RL methods and GFlowNets are able to maintain high oracle pTM scores by optimizing on the proxy pTM (Fig  4 and Table 3), while cutting down the runtime by an order of magnitude (Table  5). All results are averaged over three random seeds, and all experiments for this analysis were run on a single RTX8000 GPU. We present the complete results in Table 3.
@@ Acknowledgements
The authors would like to thank Compute Canada and Mila for the computing resources required for this work. JS would like to thank Kausthubh Ramachandran, Michael Jendrusch, Moksh Jain and Nanda Harishankar Krishna for fruitful discussions and for providing feedback on initial drafts of the paper. DN and SEK would also like to acknowledge CIFAR.
@@ 75. Guidelines:
Appendix A Additional Figures
Figure 7: Pareto plots for optimizing proteins of length $50$ on ESMFold pTM scores: Plotted across two structural diversity metrics and one sequence diversity metric (MP-HD). Only methods with atleast $0.5$ pTM are highlighted.
Appendix B Experiments on Longer Sequence Lengths
@@
Protein Confidence Diversity Score pTM ($\uparrow$) pLDDT ($\uparrow$) MP-TM ($\downarrow$) MP-RMSD ($\uparrow$) MP-HD ($\uparrow$) Random 0.40$\pm$0.00 46.9$\pm$0.48 0.35$\pm$0.01 3.99$\pm$0.02 0.90$\pm$0.02 MH-MCMC 0.93$\pm$0.00 92.2$\pm$1.09 1.00$\pm$0.00 0.04$\pm$0.03 0.01$\pm$0.01 DQN 0.37$\pm$0.03 42.2$\pm$2.74 0.62$\pm$0.14 2.76$\pm$0.75 0.15$\pm$0.12 SAC 0.71$\pm$0.02 78.5$\pm$2.32 0.71$\pm$0.12 1.87$\pm$0.43 0.66$\pm$0.06 PPO 0.86$\pm$0.02 88.9$\pm$1.90 1.00$\pm$0.00 0.03$\pm$0.03 0.01$\pm$0.01 PPO-RND 0.50$\pm$0.05 55.4$\pm$5.16 0.91$\pm$0.11 0.53$\pm$0.64 0.16$\pm$0.20 Rainbow 0.45$\pm$0.06 52.7$\pm$3.70 1.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 GFN $(\beta=6)$ 0.42$\pm$0.03 48.6$\pm$1.13 0.35$\pm$0.01 3.99$\pm$0.11 0.90$\pm$0.01 GFN $(\beta=32)$ 0.46$\pm$0.01 53.7$\pm$1.39 0.38$\pm$0.01 3.59$\pm$0.10 0.88$\pm$0.02 GFN $(\beta=64)$ 0.55$\pm$0.04 63.6$\pm$1.87 0.55$\pm$0.03 2.66$\pm$0.33 0.78$\pm$0.07
Table 5: Optimizing on pTM scores from ESMFold (Sequence length $150$): Final scores of the generated protein sequences with a budget of $10\times 10^{7}$ oracle reward evaluations. Best protein confidence scores and diversity scores are highlighted. All scores are computed on the outputs of the 3B ESMFold model.
Protein Confidence Diversity Score pTM ($\uparrow$) pLDDT ($\uparrow$) MP-TM ($\downarrow$) MP-RMSD ($\uparrow$) MP-HD ($\uparrow$) MH-MCMC 0.93$\pm$0.00 89.77$\pm$0.60 1.00$\pm$0.00 0.03$\pm$0.02 0.01$\pm$0.00 SAC 0.36$\pm$0.02 38.22$\pm$0.93 0.34$\pm$0.01 4.81$\pm$0.12 0.87$\pm$0.02 PPO 0.91$\pm$0.01 87.64$\pm$1.55 1.00$\pm$0.00 0.11$\pm$0.04 0.03$\pm$0.01 PPO-RND 0.39$\pm$0.02 43.65$\pm$5.36 0.83$\pm$0.01 1.24$\pm$0.18 0.24$\pm$0.03 GFN $(\beta=6)$ 0.35$\pm$0.01 37.70$\pm$1.96 0.34$\pm$0.02 4.74$\pm$0.21 0.86$\pm$0.03 GFN $(\beta=32)$ 0.36$\pm$0.01 39.97$\pm$2.23 0.35$\pm$0.03 4.67$\pm$0.25 0.87$\pm$0.02 GFN $(\beta=64)$ 0.40$\pm$0.03 42.04$\pm$1.07 0.62$\pm$0.01 2.62$\pm$2.15 0.52$\pm$0.46
Appendix C Biophysical Measures
In this section, we present several biophysical properties computed from the amino acid sequences of the generated proteins using Biopython [Cock et al., 2009]. Namely, we report the molecular weight ($W_{mol}$), instability index, isoelectric point, grand average of hydropathy (GRAVY) computed with the Kyte-Doolittle scale Kyte and Doolittle [1982]. These properties are then compared to a reference distribution defined by naturally occurring proteins from the Protein Data Bank (PDB) [Burley et al., 2018] using the Distributional Conformity Score (DCS) introduced in Frey et al. [2023]. Tables 6, 7, and 8 report the metrics for proteins of residue lengths $50,100$ and $150$ respectively.
Table 6: Biophysical metrics of $50$-residue proteins generated by optimizing pTM scores using ESMFold. The last column is the Distributional Conformity Score comparing generated proteins to reference proteins of the same length in PDB.
$W_{mol}$ Instability Index Isoelectric Point (pI) GRAVY DCS ($\uparrow$) MH-MCMC $5598.95$ $35.90$ $5.87$ $0.22$ $0.53$ DQN $5990.59$ $41.57$ $7.69$ $-0.16$ $0.31$ SAC $6066.13$ $47.09$ $7.75$ $0.30$ $0.39$ PPO $5955.30$ $49.38$ $6.33$ $0.57$ $0.56$ PPO-RND $6078.18$ $54.13$ $8.11$ $0.18$ $0.41$ Rainbow $5908.24$ $41.17$ $6.61$ $0.37$ $0.47$ GFN $(\beta=6)$ $6166.45$ $44.34$ $8.04$ $-0.36$ $0.29$ GFN $(\beta=32)$ $6106.51$ $51.47$ $9.18$ $-0.08$ $0.41$ GFN $(\beta=64)$ $6071.02$ $54.17$ $8.25$ $-0.29$ $0.41$
Table 7: Biophysical metrics of $100$-residue proteins generated by optimizing pTM scores using ESMFold. The last column is the Distributional Conformity Score comparing generated proteins to reference proteins of the same length in PDB.
$W_{mol}$ Instability Index Isoelectric Point (pI) GRAVY DCS ($\uparrow$) MH-MCMC $11470.28$ $35.67$ $5.52$ $-0.21$ $0.53$ DQN $12096.87$ $43.84$ $8.74$ $-0.54$ $0.29$ SAC $12079.58$ $46.09$ $6.78$ $0.87$ $0.27$ PPO $11697.12$ $34.69$ $10.55$ $0.36$ $0.52$ PPO-RND $12009.17$ $51.34$ $7.10$ $-0.52$ $0.33$ Rainbow $12121.63$ $51.81$ $9.25$ $-0.33$ $0.17$ GFN $(\beta=6)$ $12058.39$ $44.49$ $7.55$ $-0.46$ $0.34$ GFN $(\beta=32)$ $12091.40$ $44.00$ $7.47$ $-0.35$ $0.29$ GFN $(\beta=64)$ $12228.52$ $45.64$ $8.23$ $-0.14$ $0.18$
Table 8: Biophysical metrics of $150$-residue proteins generated by optimizing pTM scores using ESMFold. The last column is the Distributional Conformity Score comparing generated proteins to reference proteins of the same length in PDB.
$W_{mol}$ Instability Index Isoelectric Point (pI) GRAVY DCS ($\uparrow$) MH-MCMC $17818.09$ $32.19$ $7.99$ $-0.16$ $0.09$ SAC $17841.44$ $41.99$ $7.17$ $-0.47$ $0.22$ PPO $17971.68$ $48.31$ $7.53$ $0.09$ $0.03$ PPO-RND $17920.13$ $50.46$ $8.87$ $-0.37$ $0.19$ GFN $(\beta=6)$ $18031.62$ $41.95$ $7.15$ $-0.45$ $0.16$ GFN $(\beta=32)$ $18008.17$ $46.20$ $7.49$ $-0.39$ $0.15$ GFN $(\beta=64)$ $18129.90$ $38.15$ $7.19$ $-0.53$ $0.09$
Table 9: Optimizing on pTM scores from finetuned proxy model (ESM-PF): Distributional Conformity Score
DCS $\uparrow$
$L=50$ $L=100$
PPO 0.3074 0.1163
SAC 0.3361 0.3280
DQN 0.4807 0.1397
PPO-RND 0.3268 0.0457
Rainbow 0.4340 0.6140
GFN $(\beta=6)$ 0.3622 0.2395
GFN $(\beta=32)$ 0.3667 0.3803
GFN $(\beta=64)$ 0.2908 0.2376
MH-MCMC 0.2583 0.0321
Table 10: Optimizing on pTM scores from ESMFold: Mean absolute error in normalized frequency of amino acids between proposed sequences and naturally occurring proteins from the PDB.
MAE $\downarrow$
$L=50$ $L=100$ $L=150$
PPO 0.047 0.029 0.025
SAC 0.036 0.036 0.021
DQN 0.027 0.022 0.019
PPO-RND 0.032 0.017 0.021
Rainbow 0.024 0.018 0.050
GFN $(\beta=6)$ 0.021 0.016 0.023
GFN $(\beta=32)$ 0.025 0.018 0.022
GFN $(\beta=64)$ 0.026 0.021 0.021
MH-MCMC 0.031 0.017 0.023
Table 11: Optimizing on pTM scores from finetuned proxy model (ESM-PF): Mean absolute error in normalized frequency of amino acids between proposed sequences and naturally occurring proteins from PDB
MAE $\downarrow$
$L=50$ $L=100$
PPO 0.0526 0.0285
SAC 0.0462 0.0390
DQN 0.0207 0.0230
PPO-RND 0.0307 0.0204
Rainbow 0.0321 0.0265
GFN $(\beta=6)$ 0.0228 0.0195
GFN $(\beta=32)$ 0.0275 0.0233
GFN $(\beta=64)$ 0.0363 0.0212
MH-MCMC 0.0203 0.0240